# Fuzzy dark matter simulations

### v1

@@ 1.1.2 Fluid formulation
$\frac{\partial\rho\bm{v}}{\partial t}+\bm{\nabla}\cdot\left(\rho\bm{v}\bm{v}+\bm{\Sigma}\right)+\rho\bm{\nabla}V=0,$ (17)
@@
$\Sigma_{ij}=\frac{\hbar^{2}}{4m^{2}}\left(\frac{1}{\rho}\frac{\partial\rho}{\partial x_{i}}\frac{\partial\rho}{\partial x_{j}}-\delta_{ij}\nabla^{2}\rho\right)\quad\text{or}\quad-\frac{\hbar^{2}}{4m^{2}}\left(\rho\frac{\partial^{2}\log\rho}{\partial x_{i}\partial x_{j}}\right),$ (18)
@@ 1.2.1 Suppression of small-scale structure
Similar to Eq. (20), the suppression term $1+(M_{\rm h}/0.42M_{1/2})^{-1.1}]^{-2.2}$ is redshift-independent, as $k_{\rm J}>k_{\rm J,eq}$ throughout the matter-dominated era. Furthermore, although Eq. (22) was originally derived from collisionless $N$-body simulations with FDM initial conditions (see Section 2.6), it has been confirmed to agree well with genuine FDM wave simulations solving Eqs. (45) (May and Springel 2022). This agreement demonstrates that the FDM halo mass function in cosmological simulations is governed primarily by the initial conditions, with dynamical effects of quantum pressure playing a subdominant role.
@@
Figure 1: Projected density distributions from three cosmological simulations at $z=0$: (left) a genuine CDM simulation, (middle) a collisionless $N$-body simulation from an FDM initial condition (IC) with $m_{22}=0.1$, and (right) a genuine FDM simulation solving the Schrödinger–Poisson equations and their fluid formulation using the same FDM initial condition as the middle panel. Low-mass halos in the latter two cases are significantly suppressed in a very consistent manner, suggesting that the initial conditions dominate over the dynamical effects of quantum pressure in determining the FDM halo mass function
@@
Fig. 1 illustrates these features by comparing the density distributions at $z=0$ from three distinct simulations: collisionless $N$-body simulations initialized with CDM and FDM ($m_{22}=0.1$) power spectra, and a genuine FDM simulation using a hybrid fluid–wave algorithm (see Section 2.3) with the same FDM initial condition. The simulation box size is $5.9\,\rm Mpc$. The genuine CDM simulation exhibits significantly more low-mass halos. By contrast, the halo distributions in the latter two simulations are similar, although only the genuine FDM simulation can resolve small-scale interference fringes and solitons (see Section 1.2.2 and Section 1.2.3). The two surviving massive halos have a mass of $M_{\rm h}=4.5\times 10^{11}\,\rm{M}_{\odot}$ and $9.1\times 10^{11}\,\rm{M}_{\odot}$, respectively, close to the half-mode mass $M_{1/2}\approx 8.2\times 10^{11}\,\rm{M}_{\odot}$. The $N$-body and FDM simulations are performed using the codes GADGET-2 (Springel 2005) and GAMER (Schive et al. 2018), respectively.
@@
The suppression of small-scale structure provides a powerful means to constrain $m_{22}$. Examples include the Lyman-alpha forest power spectrum (Iršič et al. 2017; Armengaud et al. 2017; Kobayashi et al. 2017; Leong et al. 2019; Nori et al. 2019; Rogers and Peiris 2021), the cosmic microwave background (Hložek et al. 2015, 2018), the subhalo mass function (Du et al. 2017a; Benito et al. 2020; Schutz 2020; Banik et al. 2021; Nadler et al. 2021, 2025), the halo abundance and the faint end of the luminosity function at high redshifts (Bozek et al. 2015; Schive et al. 2016; Corasaniti et al. 2017; Menci et al. 2017; Schive and Chiueh 2018; Ni et al. 2019; Winch et al. 2024; Sipple et al. 2025), and the 21-cm signal during cosmic dawn (Hotinli et al. 2022).
@@ 1.2.2 Solitons
which is redshift-independent. See Fig. 12 for an illustration. Here, $r_{\rm s}$ is the soliton radius defined by $\rho_{\rm s}(r_{\rm s})=\rho_{\rm s}(0)/2$. The corresponding soliton mass $M_{\rm s}$ within $r_{\rm s}$ is
@@
Soliton properties can be inferred from their host halos, known as the soliton–halo relation. Its exact form remains under debate, with proposed scalings including $M_{\rm s}\propto M_{\rm h}^{1/3}$ (Schive et al. 2014b; Du et al. 2017b; Veltmaat et al. 2018; Chavanis 2019) and $M_{\rm s}\propto M_{\rm h}^{5/9}$ (Mocz et al. 2017; Nori and Baldi 2021). This relation can be interpreted as a manifestation of thermal equilibrium between the soliton and its surrounding halo granules (Liao et al. 2025). The transition from soliton to halo occurs at $r\approx 3.3\text{--}3.5\,r_{\rm s}$ (Mocz et al. 2017; Chiang et al. 2021), beyond which the average radial density profile follows the NFW model and closely matches the result from a collisionless $N$-body simulation using the same FDM initial condition (Liao et al. 2025, see also the middle and right panels in Fig. 1). In contrast to CDM, the concentration parameter of FDM halos decreases for $M_{\rm h}\lesssim M_{1/2}$, as a consequence of the delayed onset of halo formation (Laroche et al. 2022; Kawaii et al. 2024).
The flat core and mass excess associated with the central soliton is a distinct prediction of FDM, setting it apart from other dark matter scenarios and providing a stringent test of the model. For instance, one can constrain the soliton profile from the size and kinematic data of dwarf galaxies (Schive et al. 2014a; Calabrese and Spergel 2016; Chen et al. 2017; González-Morales et al. 2017; Burkert 2020; Hayashi et al. 2021; Zimmermann et al. 2025), ultra-diffuse galaxies (Wasserman et al. 2019; Montes et al. 2024), and the central molecular zone of the Milky Way (Li et al. 2020). The deep gravitational potential of a dense compact soliton can (i) accelerate gas accretion and thereby boost the growth of supermassive black holes in the early universe (Chiu et al. 2025), and (ii) produce a characteristic peak in galaxy rotation curves at $r\approx 1.96\,r_{\rm s}$ with a peak velocity $v_{\rm max}\approx 36\,m_{22}(M_{\rm s}/10^{8}\,\rm{M}_{\odot})\,\rm km\,\rm s^{-1}$ (Bernal et al. 2018; Bar et al. 2018, 2019; Crǎciun and Harko 2020; Bar et al. 2022; Khelashvili et al. 2023). This additional potential can also counteract tidal forces, making FDM subhalos more resilient to tidal disruption compared to their CDM counterparts (Chan et al. 2025). On the other hand, if the tidal radius is comparable to $r_{\rm s}$, solitons become unstable and undergo runaway disruption (Du et al. 2018). Furthermore, solitons exhibit density oscillations and random walk due to wave interference (Veltmaat et al. 2018; Schive et al. 2020; Li et al. 2021; Chiang et al. 2021; Dutta Chowdhury et al. 2021; Zagorac et al. 2022; Chiueh and Hsu 2023), which may dynamically heat or even tidally disrupt nearby stellar systems (Marsh and Niemeyer 2019; Schive et al. 2020; Dutta Chowdhury et al. 2023; Teodori et al. 2025).
@@ 1.2.3 Density granulation
The dynamical effects of these density fluctuations can be modeled by treating FDM granules as quasiparticles (Bar-Or et al. 2019; El-Zant et al. 2020; Chavanis 2021; Dutta Chowdhury et al. 2021). Assuming a Maxwellian velocity distribution, the effective particle mass is
@@
The perturbed gravitational field induced by FDM density granulation can scatter stars and gas, leaving observable imprints that may help distinguish FDM from other dark matter models. Examples include the outward diffusion of nuclear objects (Dutta Chowdhury et al. 2021) and stellar streams (Dalal et al. 2021), as well as the dynamical heating of star clusters (Marsh and Niemeyer 2019), dwarf galaxies (Dalal and Kravtsov 2022; Dutta Chowdhury et al. 2023; Teodori et al. 2025; Yang et al. 2025b, c), and galactic disks (Church et al. 2019; Chiang et al. 2023; Yang et al. 2024). These interference effects can also influence dynamical friction (Lancaster et al. 2020; Wang and Easther 2022; Vicente and Cardoso 2022; Vitsos and Gourgouliatos 2023; Foote et al. 2023), gravitational lensing (Chan et al. 2020; Laroche et al. 2022; Amruth et al. 2023; Powell et al. 2023), and pulsar timing signals (Khmelnitsky and Rubakov 2014; De Martino et al. 2017; Blas et al. 2017; Smarra et al. 2023). See also Hui et al. (2017); Bar-Or et al. (2019); El-Zant et al. (2020); Chavanis (2021) for further discussions.
@@ 2 Numerical methods
This section introduces a variety of FDM algorithms: wave-based schemes (Section 2.1), fluid-based schemes (Section 2.2), hybrid approaches (Section 2.3), adaptive mesh refinement (Section 2.4), eigenmode methods (Section 2.5), and collisionless $N$-body methods (Section 2.6).
@@ 2.1 Wave-based methods
The time integration of the Schrödinger–Poisson equations can be approximated to second-order accuracy using a split-step method (Taha and Ablowitz 1984; Woo and Chiueh 2009; Edwards et al. 2018; Angulo and Hahn 2022):
@@
represent the ‘kick’ and ‘drift’ operators, respectively. This method separates the nonlinear evolution due to the potential term (kick) from the linear evolution due to the kinetic term (drift). For a given wave function and potential at time $t$, the evolution proceeds by first applying the kick operator for a half time step, followed by the drift operator for a full time step. The potential is then recomputed at $t+\Delta t$, and a final half-step kick is applied. This procedure is analogous to the ‘kick-drift-kick (KDK)’ scheme commonly employed in $N$-body simulations. To enhance computational efficiency, the initial kick of the current step and the last kick of the previous step can be combined into a single full-step update. Higher-order schemes are also available (Levkov et al. 2018; Schwabe et al. 2020).
The kick operator involves only algebraic operations in position space. Specifically, it rotates the phase of the wave function while leaving the density unchanged. The corresponding time-step constraint can thus be determined by limiting the phase change per time step to below $2\pi$, thereby avoiding phase aliasing:
@@
where $\eta_{\mathcal{K}}\leq 1$ is a safety factor, and $\|V\|_{\rm max}$ denotes the maximum absolute value of the potential.
@@ 2.1.3 Local pseudo-spectral methods
In Step (a), ghost-cell data of size $N_{\rm ghost}$ are added to each side of a target grid of size $N$. These ghost cells can be filled by applying physical boundary conditions, copying from adjacent grids at the same resolution, or, in AMR simulations, interpolating from nearby coarse grids. This yields a physical domain of size $N+2N_{\rm ghost}$. Importantly, ghost cells serve a dual purpose. In addition to enabling data exchange among neighboring grids, they also help prevent unphysical data in the periodically extended region from contaminating the target grid, provided that an appropriate time-step constraint is imposed.
@@
where the filter parameters are set to $\mu=16\log(10)$, $\nu=50$, and $k_{\rm max}=\pi/\Delta h$ by default (Albin and Bruno 2011). This combination of Taylor truncation and exponential filtering ensures numerical stability by suppressing spurious high-$k$ modes.
Finally, in Step (d), only the data within the original grid of size $N$ are retained.
@@ 2.2.1 Smoothed particle hydrodynamics
SPH offers several benefits owing to its Lagrangian nature. The method supports smoothly adaptive resolution and can be easily coupled to $N$-body gravity solvers. It is Galilean invariant, which minimizes advection errors relative to Eulerian methods. However, it faces several significant challenges. The quantum potential $Q$ depends on higher-order spatial derivatives of the density, which can be very noisy with standard SPH derivative estimators. Both $Q$ and velocity diverge at vortices where $\rho\rightarrow 0$ (see Section 3.2 and Fig. 14), which can cause substantial truncation errors, hinder numerical convergence, and force prohibitively small time steps. The smoothing length, generally scaling as $h\propto\rho^{-1/3}$ (for fixed particle mass), can exceed the local de Broglie wavelength $\lambda_{\rm dB}$ in low-density regions, making it difficult to resolve density granulation and the associated quantum pressure within halos.
These limitations likely contribute to the following discrepancies between different SPH implementations and wave-based methods. First, it remains uncertain whether SPH can robustly resolve solitons and recover the soliton–halo relation in FDM halos. For example, the flat cores found in the SPH spherical-collapse and cosmological simulations of Zhang et al. (2018b, Fig. 4), Nori and Baldi (2021, Fig. 4), and Nori et al. (2023, Fig. 7) do not exhibit the sharp soliton–halo transition commonly predicted in wave-based simulations (e.g., see Fig. 13) and observed in the SPH soliton-collision simulations of Veltmaat and Niemeyer (2016, Fig. 4). Relatedly, the velocity dispersion within the flat cores in Zhang et al. (2018b, Fig. 5) and Nori et al. (2023, Fig. 8) remains roughly constant. This is inconsistent with the soliton solution, which is supported by quantum pressure rather than velocity dispersion, as confirmed by wave-based simulations (e.g., Liao et al. 2025). Furthermore, Nori and Baldi (2018) found that the matter power spectrum is suppressed on small scales by the quantum pressure relative to $N$-body simulations. This contrasts with the results of the SPH simulations by Veltmaat and Niemeyer (2016) and the wave-based global spectral simulations by May and Springel (2021), which reported enhanced small-scale power at low redshifts due to wave interference. These limitations and controversies highlight ongoing uncertainties and raise concerns about the reliability of SPH simulations in capturing the coarse-grained properties of FDM.

### v2

@@ 1.1.2 Fluid formulation
$\frac{\partial\rho\bm{v}}{\partial t}+\bm{\nabla}\cdot\left(\rho\bm{v}\otimes\bm{v}+\bm{\Sigma}\right)+\rho\bm{\nabla}V=0,$ (17)
@@
$\Sigma_{ij}=\frac{\hbar^{2}}{4m^{2}}\left(\frac{1}{\rho}\frac{\partial\rho}{\partial x_{i}}\frac{\partial\rho}{\partial x_{j}}-\delta_{ij}\nabla^{2}\rho\right)\quad\text{or}\quad-\frac{\hbar^{2}}{4m^{2}}\left(\rho\frac{\partial^{2}\ln\rho}{\partial x_{i}\partial x_{j}}\right),$ (18)
@@ 1.2.1 Suppression of small-scale structure
Similar to Eq. (20), the suppression term $1+(M_{\rm h}/0.42M_{1/2})^{-1.1}]^{-2.2}$ is redshift-independent, as $k_{\rm J}>k_{\rm J,eq}$ throughout the matter-dominated era. Furthermore, although Eq. (22) was originally derived from collisionless NN-body simulations with FDM initial conditions (see Section 2.6), it has been confirmed to agree well with genuine FDM wave simulations solving Eqs. (45) (May and Springel 2022). This agreement demonstrates that the FDM halo mass function in cosmological simulations is governed primarily by the initial conditions, with dynamical effects of quantum pressure playing a subdominant role.
@@
Figure 1: Projected density distributions from three cosmological simulations at $z=0$: (left) a genuine CDM simulation, (middle) a collisionless NN-body simulation from an FDM initial condition (IC) with $m_{22}=0.1$, and (right) a genuine FDM simulation solving the Schrödinger–Poisson equations and their fluid formulation using the same FDM initial condition as the middle panel. Low-mass halos in the latter two cases are significantly suppressed in a very consistent manner, suggesting that the initial conditions dominate over the dynamical effects of quantum pressure in determining the FDM halo mass function
@@
Fig. 1 illustrates these features by comparing the density distributions at $z=0$ from three distinct simulations: collisionless NN-body simulations initialized with CDM and FDM ($m_{22}=0.1$) power spectra, and a genuine FDM simulation using a hybrid fluid–wave algorithm (see Section 2.3) with the same FDM initial condition. The simulation box size is $5.9\,\rm Mpc$. The genuine CDM simulation exhibits significantly more low-mass halos. By contrast, the halo distributions in the latter two simulations are similar, although only the genuine FDM simulation can resolve small-scale interference fringes and solitons (see Section 1.2.2 and Section 1.2.3). The two surviving massive halos have a mass of $M_{\rm h}=4.5\times 10^{11}\,\rm{M}_{\odot}$ and $9.1\times 10^{11}\,\rm{M}_{\odot}$, respectively, close to the half-mode mass $M_{1/2}\approx 8.2\times 10^{11}\,\rm{M}_{\odot}$. The NN-body and FDM simulations are performed using the codes GADGET-2 (Springel 2005) and GAMER (Schive et al. 2018), respectively.
@@
The suppression of small-scale structure provides a powerful means to constrain $m_{22}$. Examples include the Lyman-alpha forest power spectrum (Iršič et al. 2017; Armengaud et al. 2017; Kobayashi et al. 2017; Leong et al. 2019; Nori et al. 2019; Rogers and Peiris 2021), the cosmic microwave background (Hložek et al. 2015, 2018), the subhalo mass function (Du et al. 2017a; Benito et al. 2020; Schutz 2020; Banik et al. 2021; Nadler et al. 2021, 2025), the halo abundance and the faint end of the luminosity function at high redshifts (Bozek et al. 2015; Schive et al. 2016; Corasaniti et al. 2017; Menci et al. 2017; Schive and Chiueh 2018; Ni et al. 2019; Winch et al. 2024; Sipple et al. 2025), the galaxy shear weak lensing power spectrum (Dentler et al. 2022), and the 21-cm signal during cosmic dawn (Hotinli et al. 2022).
@@ 1.2.2 Solitons
which is redshift-independent as the soliton is an object detached from cosmological expansion. See Fig. 12 for an illustration. Here, $r_{\rm s}$ is the soliton radius defined by $\rho_{\rm s}(r_{\rm s})=\rho_{\rm s}(0)/2$. The corresponding soliton mass $M_{\rm s}$ within $r_{\rm s}$ is
@@
Soliton properties can be inferred from their host halos, known as the soliton–halo relation. Its exact form remains under debate, with proposed scalings including $M_{\rm s}\propto M_{\rm h}^{1/3}$ (Schive et al. 2014b; Du et al. 2017b; Veltmaat et al. 2018; Chavanis 2019) and $M_{\rm s}\propto M_{\rm h}^{5/9}$ (Mocz et al. 2017; Nori and Baldi 2021). This relation can be interpreted as a manifestation of thermal equilibrium between the soliton and its surrounding halo granules (Liao et al. 2025). The transition from soliton to halo occurs at $r\approx 3.3\text{--}3.5\,r_{\rm s}$ (Mocz et al. 2017; Chiang et al. 2021), beyond which the average radial density profile follows the NFW model and closely matches the result from a collisionless NN-body simulation using the same FDM initial condition (Liao et al. 2025, see also the middle and right panels in Fig. 1). In contrast to CDM, the concentration parameter of FDM halos decreases for $M_{\rm h}\lesssim M_{1/2}$, as a consequence of the delayed onset of halo formation (Laroche et al. 2022; Kawaii et al. 2024).
The flat core and mass excess associated with the central soliton is a distinct prediction of FDM, setting it apart from other dark matter scenarios and providing a stringent test of the model. For instance, one can constrain the soliton profile from the size and kinematic data of dwarf galaxies (Schive et al. 2014a; Calabrese and Spergel 2016; Chen et al. 2017; González-Morales et al. 2017; Burkert 2020; Safarzadeh and Spergel 2020; Hayashi et al. 2021; Zimmermann et al. 2025), ultra-diffuse galaxies (Wasserman et al. 2019; Montes et al. 2024), and the central molecular zone of the Milky Way (Li et al. 2020). The deep gravitational potential of a dense compact soliton can (i) accelerate gas accretion and thereby boost the growth of supermassive black holes in the early universe (Chiu et al. 2025), and (ii) produce a characteristic peak in galaxy rotation curves at $r\approx 1.96\,r_{\rm s}$ with a peak velocity $v_{\rm max}\approx 36\,m_{22}(M_{\rm s}/10^{8}\,\rm{M}_{\odot})\,\rm km\,\rm s^{-1}$ (Bernal et al. 2018; Bar et al. 2018, 2019; Crǎciun and Harko 2020; Chan and Fai Yeung 2021; Bar et al. 2022; Khelashvili et al. 2023). This additional potential can also counteract tidal forces, making FDM subhalos more resilient to tidal disruption compared to their CDM counterparts (Chan et al. 2025). On the other hand, if the tidal radius is comparable to $r_{\rm s}$, solitons become unstable and undergo runaway disruption (Du et al. 2018). Furthermore, solitons exhibit density oscillations and random walk due to wave interference (Veltmaat et al. 2018; Schive et al. 2020; Li et al. 2021; Chiang et al. 2021; Dutta Chowdhury et al. 2021; Zagorac et al. 2022; Chiueh and Hsu 2023), which may dynamically heat or even tidally disrupt nearby stellar systems (Marsh and Niemeyer 2019; Schive et al. 2020; Dutta Chowdhury et al. 2023; Teodori et al. 2026).
@@ 1.2.3 Density granulation
The dynamical effects of these density fluctuations can be modeled by treating FDM granules as quasiparticles (Bar-Or et al. 2019; Church et al. 2019; El-Zant et al. 2020; Chavanis 2021; Dutta Chowdhury et al. 2021). Assuming a Maxwellian velocity distribution, the effective particle mass is
@@
The perturbed gravitational field induced by FDM density granulation can scatter stars and gas, leaving observable imprints that may help distinguish FDM from other dark matter models. Examples include the outward diffusion of nuclear objects (Dutta Chowdhury et al. 2021) and stellar streams (Dalal et al. 2021), as well as the dynamical heating of star clusters (Marsh and Niemeyer 2019), dwarf galaxies (Dalal and Kravtsov 2022; Dutta Chowdhury et al. 2023; Teodori et al. 2026; Yang et al. 2025b, c), and galactic disks (Church et al. 2019; Chiang et al. 2023; Yang et al. 2024). These interference effects can also influence dynamical friction (Lancaster et al. 2020; Wang and Easther 2022; Vicente and Cardoso 2022; Vitsos and Gourgouliatos 2023; Foote et al. 2023), gravitational lensing (Chan et al. 2020; Laroche et al. 2022; Amruth et al. 2023; Powell et al. 2023), and pulsar timing signals (Khmelnitsky and Rubakov 2014; De Martino et al. 2017; Blas et al. 2017; Smarra et al. 2023). See also Hui et al. (2017); Bar-Or et al. (2019); El-Zant et al. (2020); Chavanis (2021) for further discussions.
@@ 2 Numerical methods
This section introduces a variety of FDM algorithms: wave-based schemes (Section 2.1), fluid-based schemes (Section 2.2), hybrid approaches (Section 2.3), adaptive mesh refinement (Section 2.4), eigenmode methods (Section 2.5), and collisionless NN-body methods (Section 2.6).
@@ 2.1 Wave-based methods
The time integration of the Schrödinger–Poisson equations can be approximated to second-order accuracy using a split-step method (Taha and Ablowitz 1984; Woo and Chiueh 2009; Mocz et al. 2017; Edwards et al. 2018; Angulo and Hahn 2022):
@@
represent the ‘kick’ and ‘drift’ operators, respectively. This method separates the nonlinear evolution due to the potential term (kick) from the linear evolution due to the kinetic term (drift). For a given wave function and potential at time $t$, the evolution proceeds by first applying the kick operator for a half time step, followed by the drift operator for a full time step. The potential is then recomputed at $t+\Delta t$, and a final half-step kick is applied. This procedure is analogous to the ‘kick-drift-kick (KDK)’ scheme commonly employed in NN-body simulations. To enhance computational efficiency, the initial kick of the current step and the last kick of the previous step can be combined into a single full-step update. Higher-order schemes are also available (Levkov et al. 2018; Schwabe et al. 2020).
The kick operator involves only algebraic operations in position space. Specifically, it rotates the phase of the wave function while leaving the density unchanged. The corresponding time-step constraint is commonly set by limiting the phase change per time step to below $2\pi$ (Feit et al. 1982), thereby avoiding phase aliasing:
@@
where $\eta_{\mathcal{K}}\leq 1$ is a safety factor, and $\|V\|_{\rm max}$ denotes the maximum absolute value of the potential. However, since only phase differences are dynamically relevant, it may be more physically motivated to require that the relative phase variation between adjacent simulation elements per time step remain below $\pi$.
@@ 2.1.3 Local pseudo-spectral methods
In Step (a), ghost-cell data of size $N_{\rm ghost}$ are added to each side of a target grid of size NN. These ghost cells can be filled by applying physical boundary conditions, copying from adjacent grids at the same resolution, or, in AMR simulations, interpolating from nearby coarse grids. This yields a physical domain of size $N+2N_{\rm ghost}$. Importantly, ghost cells serve a dual purpose. In addition to enabling data exchange among neighboring grids, they also help prevent unphysical data in the periodically extended region from contaminating the target grid, provided that an appropriate time-step constraint is imposed.
@@
where the filter parameters are set to $\mu=16\ln(10)$, $\nu=50$, and $k_{\rm max}=\pi/\Delta h$ by default (Albin and Bruno 2011). This combination of Taylor truncation and exponential filtering ensures numerical stability by suppressing spurious high-$k$ modes.
Finally, in Step (d), only the data within the original grid of size NN are retained.
@@ 2.2.1 Smoothed particle hydrodynamics
SPH offers several benefits owing to its Lagrangian nature. The method supports smoothly adaptive resolution and can be easily coupled to NN-body gravity solvers. It is Galilean invariant, which minimizes advection errors relative to Eulerian methods. However, it faces several significant challenges. The quantum potential $Q$ depends on higher-order spatial derivatives of the density, which can be very noisy with standard SPH derivative estimators. Both $Q$ and velocity diverge at vortices where $\rho\rightarrow 0$ (see Section 3.2 and Fig. 14), which can cause substantial truncation errors, hinder numerical convergence, and force prohibitively small time steps. The smoothing length, generally scaling as $h\propto\rho^{-1/3}$ (for fixed particle mass), can exceed the local de Broglie wavelength $\lambda_{\rm dB}$ in low-density regions, making it difficult to resolve density granulation and the associated quantum pressure within halos.
These limitations likely contribute to the following discrepancies between different SPH implementations and wave-based methods. First, it remains uncertain whether SPH can robustly resolve solitons and recover the soliton–halo relation in FDM halos. For example, the flat cores found in the SPH spherical-collapse and cosmological simulations of Zhang et al. (2018b, Fig. 4), Nori and Baldi (2021, Fig. 4), and Nori et al. (2023, Fig. 7) do not exhibit the sharp soliton–halo transition commonly predicted in wave-based simulations (e.g., see Fig. 13) and observed in the SPH soliton-collision simulations of Veltmaat and Niemeyer (2016, Fig. 4). Relatedly, the velocity dispersion within the flat cores in Zhang et al. (2018b, Fig. 5) and Nori et al. (2023, Fig. 8) remains roughly constant. This is inconsistent with the soliton solution, which is supported by quantum pressure rather than velocity dispersion, as confirmed by wave-based simulations (e.g., Liao et al. 2025). Furthermore, Nori and Baldi (2018) found that the matter power spectrum is suppressed on small scales by the quantum pressure relative to NN-body simulations. This contrasts with the results of the SPH simulations by Veltmaat and Niemeyer (2016) and the wave-based global spectral simulations by May and Springel (2021), which reported enhanced small-scale power at low redshifts due to wave interference. These limitations and controversies highlight ongoing uncertainties and raise concerns about the reliability of SPH simulations in capturing the coarse-grained properties of FDM.